OpenStudy (kunals0424):

There are four naturally occurring isotopes of the element chromium. The relative abundance of each is 50Cr 4.345% 49.946044 amu 52Cr 83.789% 51.940508 amu 53Cr 9.501% 52.940649 amu 54Cr 2.365% 53.93880 amu Find the average atomic mass of chromium.

OpenStudy (jiteshmeghwal9):

\[\sum \frac{(\text{Atomic mass of each element * percentage abundance})}{100}\]

OpenStudy (jiteshmeghwal9):

Average atomic mass =\[\frac{50*4.345+52*83.789+53*9.501+54*2.365}{100}\]